OpenStudy (anonymous):

Please help! A rocket of mass 1.0x10^3kg is being fired to a height of 5.0 x 10^3 m. The rocket engine shuts off when the rocket reaches a height of 1.0x10^3m, and the rocket coasts up to 5.0 x 10^3 m. A) What velocity must the rocket have at the 1.0x10^3 m point to enable it to reach 5.0 x 10^3m B) What acceleration did the rocket experience while the engine was on? off? C) What force did the rocket engine exert on the rocket?

OpenStudy (anonymous):

Strategy: Part A) Assuming rocket stops at 5.0x10^3m, use \(v^2=v_0^2-2gy\) where y is the rest of travelling distance, (5.0-1.0)x10^3m. Part B) While on: The rocket must reach velocity in a from rest at a distance 1.0x10^3m. Again, use \(v^2=v_0^2-2gy\). While off: Simply \(g\). Part C) Newtons famous \(\Sigma F=ma\) where a is the one you found in Part B).

OpenStudy (anonymous):

Thank you very much! Fanned and medaled!

OpenStudy (anonymous):

I got a question what is that v in the equation after the equals? is that 2/0?

OpenStudy (anonymous):

Tell me if im right and please fill in that variable i dont know: 1)V^2=? - 2(-9.8)(4.0x10^3) 2)V^2=? - 2(-9.8)(1.0x10^3) B) 9.8m/s^2 3) (1.0x10^3) x whatever i get from 2a?

OpenStudy (anonymous):

1 V=0; ?= v you're looking for 2 v = the one found in 1; ? = 0